# Arcade Processes for Informed Martingale Interpolation

### v2

@@ Acknowledgments
A. Macrina is grateful for discussions on stochastic interpolation with Camilo A. Garcia Trillos and Jun Sekine, especially in the context of information-based modelling and stochastic filtering, during a research visit at Osaka University in March 2017 as part of a Daiwa Anglo-Japanese Foundation Small Grant, Ref: 3078/12370. Osaka University’s hospitality is acknowledged and appreciated. A. Macrina is also grateful for the support by the Fields Institute for Research in Mathematical Sciences through the award of a Fields Research Fellowship in 2022. The contents of this manuscript are solely the responsibility of the authors and do not necessarily represent the official views of the Fields Institute. G. Kassis acknowledges the UCL Department of Mathematics for a Teaching Assistantship Award. Furthermore, the authors thank L. P. Hughston for helpful comments and suggestions at an early stage of the work on arcade processes, B. Acciaio and G. Pammer for pointers to optimal transport and stretched Brwonian motion, S. Cohen, J. Guyon, F. Krach, J. Obloj, G. W. Peters, and T.-K. L. Wong for useful conversations, and J. Akahori, K. Owari and D. C. Schwarz for feedback. Attendees of the AIFMRM Research Seminar at the University of Cape Town (Aug. 2021) and participants in the SIAM Conference on Mathematics of Data Science (MDS22, Sept. 2022), the 7th International Conference Mathematics in Finance—MiF Kruger Park— (July 2023, South Africa), the Talks in Financial and Insurance Mathematics at ETH Zurich (Nov. 2023, Switzerland), the Department of Mathematics Seminar of Ritsumeikan University (Nov. 2023, Japan), and participants in the FAMiLLY Workshop at the University of Liverpool (Dec. 2023, U.K.) and the XXV Workshop on Quantitative Finance (Apr. 2024, Italy) are thanked for their comments and suggestions.

### v3

@@ Acknowledgments
A. Macrina is grateful for discussions on stochastic interpolation with Camilo A. Garcia Trillos and Jun Sekine, especially in the context of information-based modelling and stochastic filtering, during a research visit at Osaka University in March 2017 as part of a Daiwa Anglo-Japanese Foundation Small Grant, Ref: 3078/12370. Osaka University’s hospitality is acknowledged and appreciated. A. Macrina is also grateful for the support by the Fields Institute for Research in Mathematical Sciences through the awards of a Fields Research Fellowship in 2022 and an Elliott-Yui Distinguished Visitorship in 2025. The contents of this manuscript are solely the responsibility of the authors and do not necessarily represent the official views of the Fields Institute. G. Kassis acknowledges the UCL Department of Mathematics for a Teaching Assistantship Award. Furthermore, the authors thank L. P. Hughston for helpful comments and suggestions at an early stage of the work on arcade processes, B. Acciaio and G. Pammer for pointers to optimal transport and stretched Brwonian motion, S. Cohen, J. Guyon, F. Krach, G. W. Peters, and T.-K. L. Wong for useful conversations, and J. Akahori, K. Owari and D. C. Schwarz for feedback. Attendees of the AIFMRM Research Seminar at the University of Cape Town (Aug. 2021) and participants in the SIAM Conference on Mathematics of Data Science (MDS22, Sept. 2022), the 7th International Conference Mathematics in Finance—MiF Kruger Park— (July 2023, South Africa), the Talks in Financial and Insurance Mathematics at ETH Zurich (Nov. 2023, Switzerland), of the Department of Mathematics Seminar of Ritsumeikan University (Nov. 2023, Japan), and participants in the FAMiLLY Workshop at the University of Liverpool (Dec. 2023, U.K.), the XXV Workshop on Quantitative Finance (Apr. 2024, Italy), the Research Seminar of the Department of Mathematics & Statistics, University of Ottawa (Nov. 2024, Canada), the Quantitative Methods in Finance (QMF) 2024 International Conference at the University of Technology Sydney (Dec. 2024, Australia), and in the 69th Annual Meeting of the Australian Mathematical Society (Dec. 2025, Australia) are thanked for their comments and suggestions. Last but not least, we are grateful to two anonymous reviewers and an associated editor for constructive comments that have helped with improving this paper.
Appendix A Arcade processes
In this appendix, we include further examples of arcade processes along with their simulations.
Example A.48.
For $n>1$, we can generalise the anticipative Brownian bridge by taking
$f_{0}(t)=\frac{T_{1}-t}{T_{1}-T_{0}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\frac{t-T_{n-1}}{T_{n}-T_{n-1}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (A.1)
and
$f_{i}(t)=\frac{t-T_{i-1}}{T_{i}-T_{i-1}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\frac{T_{i+1}-t}{T_{i+1}-T_{i}}\mathds{1}_{(T_{i},T_{i+1}]}(t),\quad\text{for }i=1,\ldots,n-1.$ (A.2)
We call this AP the stitched Brownian AP, for it can be written as
$A_{t}^{(n)}=\left\{\begin{array}[]{ll}B_{t}-\frac{T_{1}-t}{T_{1}-T_{0}}B_{T_{0}}-\frac{t-T_{0}}{T_{1}-T_{0}}B_{T_{1}},&\mbox{if }t\in[T_{0},T_{1}),\\ \\ B_{t}-\frac{T_{2}-t}{T_{2}-T_{1}}B_{T_{1}}-\frac{t-T_{1}}{T_{2}-T_{1}}B_{T_{2}},&\mbox{if }t\in[T_{1},T_{2}),\\ \,\vdots\\ B_{t}-\frac{T_{n}-t}{T_{n}-T_{n-1}}B_{T_{n-1}}-\frac{t-T_{n-1}}{T_{n}-T_{n-1}}B_{T_{n}},&\mbox{if }t\in[T_{n-1},T_{n}].\end{array}\right$. (A.3)
Figure 7: Paths simulation of a stitched Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\mid i=0,1,\ldots,5\}$, cf., Example A.48.
Example A.49.
Another way of generalizing the anticipative Brownian bridge to obtain an AP driven by Brownian motion is by using Lagrange’s polynomial interpolation, that is,
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}\quad\text{for }i=0,\ldots,n.$ (A.4)
$A_{t}^{(n)}=B_{t}-\sum_{i=0}^{n}\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}B_{T_{i}}.$ (A.5)
Figure 8: Paths simulation of a Lagrange-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\mid i=0,1,\ldots,5\}$, cf., Example A.49.
The Lagrange APs inherit Runge’s phenomenon from their interpolation coefficients. When $n$ is large, the AP oscillates around the edges of the interval. One can control this effect by applying a transformation to the interpolating coefficients. For instance, the map $x\mapsto|x|^{2(1-|x|)}$ applied to each of the interpolating coefficients
$f_{i}(t)=\prod\limits_{k=0,k\neq i}^{n}\frac{T_{k}-t}{T_{k}-T_{i}}$
for $i=0,\ldots,n$, yields another set of interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{n}$ which no longer suffer from Runge’s phenomenon.
Figure 9: Paths simulation of a Lagrange-Brownian AP with $n=10$, using the equidistant partition $\{T_{i}=i\mid i=0,1,\ldots,10\}$, illustrating Runge’s phenomenon, cf., Example A.49.
Figure 10: Paths simulation of a Brownian AP with $n=10$ and interpolating coefficients $\tilde{f}_{0},\ldots,\tilde{f}_{10}$, using the equidistant partition $\{T_{i}=i\mid i=0,1,\ldots,10\}$, cf., Example A.49.
Example A.50.
Elliptic APs have interpolating coefficients given by
$f_{0}(t)=\sqrt{1-\left(\frac{t-T_{0}}{T_{1}-T_{0}}\right)^{2}}\mathds{1}_{[T_{0},T_{1}]}(t),\quad f_{n}(t)=\sqrt{1-\left(\frac{t-T_{n}}{T_{n}-T_{n-1}}\right)^{2}}\mathds{1}_{(T_{n-1},T_{n}]}(t),$ (A.6)
$f_{i}(t)=\sqrt{1-\left(\frac{t-T_{i}}{T_{i}-T_{i-1}}\right)^{2}}\mathds{1}_{(T_{i-1},T_{i}]}(t)+\sqrt{1-\left(\frac{t-T_{i}}{T_{i+1}-T_{i}}\right)^{2}}\mathds{1}_{(T_{i},T_{i+1}]}(t),$ (A.7)
for $i=1,\ldots,n-1$.
Figure 11: Paths simulation of an elliptic-Brownian AP with $n=5$, using the equidistant partition $\{T_{i}=2i\mid i=0,1,\ldots,5\}$, cf., Example A.50.
Appendix B Filtered arcade martingales
In this appendix, we include further examples of filtered arcade martingales.
Example B.51.
$X_{1}\mid X_{0}=\left\{\begin{array}[]{ll}\frac{3}{2}X_{0}+\frac{1}{2}&\mbox{with probability }\frac{3}{4},\\ \\ -\frac{1}{2}X_{0}-\frac{3}{2}&\mbox{with probability }\frac{1}{4}.\end{array}\right$. (B.1)
This is a martingale coupling since $\mathbb{E}[X_{1}\mid X_{0}]=X_{0}$. In fact, it can be shown that this coupling is the solution to a martingale optimal transport problem, see [29]. For any $\{T_{0},T_{1}\}$-conditionally Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}$ $={\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]}$
$=\frac{(9X_{0}+3)f^{I_{t}^{(1)}\mid X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})-(X_{0}+3)f^{I_{t}^{(1)}\mid X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}{6f^{I_{t}^{(1)}\mid X_{0},X_{1}=\frac{3}{2}X_{0}+\frac{1}{2}}(I_{t}^{(1)})+2f^{I_{t}^{(1)}\mid X_{0},X_{1}=-\frac{1}{2}X_{0}-\frac{3}{2}}(I_{t}^{(1)})}.$ (B.2)
Example B.52.
Let $X_{0}\sim\mathcal{N}(0,1)$ and $X_{1}\sim\mathcal{N}(0,2)$, where $X_{1}\mid X_{0}\sim\mathcal{N}(X_{0},1)$. For any $\{T_{0},T_{1}\}$-conditionally Markov $X$-RAP $(I_{t}^{(1)})$, we have
$M_{t}=\mathbb{E}[X_{1}\mid X_{0},I_{t}^{(1)}]=\frac{\int_{\mathbb{R}}yf^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}{\int_{\mathbb{R}}f^{I_{t}^{(1)}\mid X_{0},X_{1}=y}(I_{t}^{(1)})\mathrm{e}^{\frac{-(y-X_{0})^{2}}{2}}\,\mathrm{d}y}.$ (B.3)
Example B.53.
Let $D_{t}=tB_{t}$ where $(B_{t})_{t\geqslant 0}$ is a standard Brownian motion. Then $K_{D}(x,y)=\min(x,y)^{2}\max(x,y)$, and so $H_{1}(x)=x^{2}$, $H_{2}(x)=x$. The standard AP on $T_{0},T_{1}]$ driven by $(D_{t})$ is given by
$A_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}D_{T_{0}}-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}D_{T_{1}}.$ (B.4)
We have
$K_{A}(x,y)=\frac{\min(x,y)(\min(x,y)-T_{0})\max(x,y)(T_{1}-\max(x,y))}{T_{1}-T_{0}},$ (B.5)
$I_{t}^{(1)}=D_{t}-\frac{T_{1}t-t^{2}}{T_{1}T_{0}-T_{0}^{2}}(D_{T_{0}}-X_{0})-\frac{t^{2}-tT_{0}}{T_{1}^{2}-T_{1}T_{0}}(D_{T_{1}}-X_{1}).$ (B.6)
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]s}{T_{1}^{2}s-s^{2}T_{1}}\,\mathrm{d}W_{s}=X_{0}+\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{s}^{(1)}]}{T_{1}^{2}-sT_{1}}\,\mathrm{d}W_{s}.$ (B.7)
Example B.54.
Let $(D_{t})$ be the Ornstein-Uhlenbeck process satisfying the SDE
$D_{t}=\int_{0}^{t}\theta\left(\mu-D_{s}\right)\,\mathrm{d}s+\int_{0}^{t}\sigma\,\mathrm{d}B_{s},$ (B.8)
$A_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{0}}-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}D_{T_{1}}.$ (B.9)
$I_{t}^{(1)}=D_{t}-\frac{\mathrm{e}^{\theta(T_{1}-t)}-\mathrm{e}^{-\theta(T_{1}-t)}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{0}}-X_{0})-\frac{\mathrm{e}^{\theta(t-T_{0})}-\mathrm{e}^{-\theta(t-T_{0})}}{\mathrm{e}^{\theta(T_{1}-T_{0})}-\mathrm{e}^{-\theta(T_{1}-T_{0})}}(D_{T_{1}}-X_{1}).$ (B.10)
$M_{t}=X_{0}+\int_{T_{0}}^{t}\frac{\sigma\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]}{\frac{\sigma^{2}}{2\theta}\left(\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}\right)}\,\mathrm{d}W_{s}=X_{0}+\frac{2\theta}{\sigma}\int_{T_{0}}^{t}\frac{\mathrm{Var}[X_{1}\mid X_{0},I_{t}^{(1)}]}{\mathrm{e}^{\theta(T_{1}-s)}-\mathrm{e}^{\theta(s-T_{1})}}\,\mathrm{d}W_{s}.$ (B.11)